OpenStudy (anonymous):

A car goes around a curve on a road that is banked at an angle of 34.5∘ . Even though the road is slick, the car will stay on the road without any friction between its tires and the road when its speed is 23.0m/s . What is the radius of the curve?

OpenStudy (anonymous):

whats the answer lol

OpenStudy (anonymous):

@Pompeii00

OpenStudy (anonymous):

running short on time.. this online hw due at 12

OpenStudy (anonymous):

its wrong. thanks thoug

OpenStudy (anonymous):

still wong

OpenStudy (anonymous):

**wrong

OpenStudy (anonymous):

@Pompeii00

OpenStudy (anonymous):

I must be tired, i just made some dumb mistakes. I apologize since you were trying to get this done in a timely manner. So let me just show a solution. This should be correct.  From this, we see that Fn cos(a) = mg Fn = mg/cos(a) Now, we are only concerned with horizontal forces, since that provides the centripetal force. Fn sin(a) = mg/cos(a) * sin(a) = mg tan(a) Centripetal force is: Fc = mv^2 / r = mg tan(a) v^2 / r = g tan(a) r = v^2 / g tan(a) = (23)^2 / (9.8)(tan(34.5)) r = 78.5 m